OpenStudy (anonymous):

Write using exponents. (–4)(–4)

OpenStudy (lgbasallote):

-4^2

sam (.sam.):

Multiply -4 by -4 to get 16. 16

OpenStudy (ajprincess):



OpenStudy (anonymous):

\[{\left(-4\right)}^2=16\]Be careful of writing -4^2 instead of (-4)^2 in HTML; most processors will evaluate -4^2=-16 while (-4)^2=16.